A lease has a term of 8 years with annual payments of $21,400. The asset would cost $131,000 to buy and would be depreciated straight-line to a zero salvage value over 8 years. The actual salvage value is zero. If the firm has a tax rate of 30 percent, what is the incremental cash flow in Year 8 of leasing rather than purchasing?

$14,856.50

-$16,375.00

-$21,400.00

$16,375.00

-$19,892.50
Answer #1

Annual lease payments = 21400

Tax rate = 30%

Net Annual lease payment = 21400*(1-0.30) = 14980

Annual depreciation = 131000/8 = 16375

Tax savings on depreciation = 16375*30% = 4912.5

Incremental cash flow of purchasing in year 8= Tax flows on depreciation - Annual cash flow in purchasing

= -4912.5 - 14980 = -19892.5